OpenStudy (anonymous):

please help. For an angle Θ with the point (-5, -12) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

tan theta = (y/x) theta=atan(y/x)